# prompt[Prompt][List of Prompts]

### v3

@@ Abstract
Recent advancements in text-to-video models such as Sora, Gen-3, MovieGen, and CogVideoX are pushing the boundaries of synthetic video generation, with adoption seen in fields like robotics, autonomous driving, and entertainment. As these models become prevalent, various metrics and benchmarks have emerged to evaluate the quality of the generated videos. However, these metrics emphasize visual quality and smoothness, neglecting temporal fidelity and text-to-video alignment, which are crucial for safety-critical applications. To address this gap, we introduce NeuS-V, a novel synthetic video evaluation metric that rigorously assesses text-to-video alignment using neuro-symbolic formal verification techniques. Our approach first converts the prompt into a formally defined Temporal Logic (TL) specification and translates the generated video into an automaton representation. Then, it evaluates the text-to-video alignment by formally checking the video automaton against the TL specification. Furthermore, we present a dataset of temporally extended prompts to evaluate state-of-the-art video generation models against our benchmark. We find that NeuS-V demonstrates a higher correlation by over 5×5$\times$5 × with human evaluations when compared to existing metrics. Our evaluation further reveals that current video generation models perform poorly on these temporally complex prompts, highlighting the need for future work in improving text-to-video generation capabilities.
@@ 1 Introduction
Various evaluation metrics have emerged to identify well-aligned synthetic videos to the prompt from recently developed T2V models [20, 14, 4, 55]. While many video evaluation tools [39, 37, 23] focus on visual quality, some approaches use vision-language models [48, 49] or video captioning models [9] to evaluate the semantic alignment between the generated video and the original prompt. More notably, VBench [23] evaluates videos across multiple dimensions and categories and is recently being used in leaderboards***https://huggingface.co/spaces/Vchitect/VBench_Leaderboard to evaluate T2V models. However, naive evaluation of a video generated by one neural network using another is neither rigorous nor interpretable. Moreover, current evaluation metrics still miss critical spatio-temporal aspects that are intrinsic to video data (See Fig. 1).
@@ 3 Preliminaries
In the following sections, we present a running example to help illustrate our approach. Let’s say we generate an autonomous driving video using T2V models with the prompt – “A car drives down a road on a clear sunny day, interacting with cyclists who signal turns to avoid obstacles”.
@@
$\begin{array}[]{rl}\mathcal{P}&=\{\text{car driving},\ \text{clear day},\ \text{cyclist signals turn},\\ &\quad\ \ \text{cyclist turns},\ \text{cyclist avoids obstacle}\},\\ \Phi&=\Box\Big{(}(\text{car driving}\wedge\text{clear day})\wedge(\text{cyclist signals turn})\\ &\quad\Rightarrow\diamondsuit(\text{cyclist turns}\wedge\text{cyclist avoids obstacle})\Big{)}.\end{array}$ (1)
@@ 4.1 Prompt Understanding via Temporal Logic Specification
We propose Prompt Understanding via temporal Logic (PULS), a novel method that facilitates an efficient and accurate conversion of text prompts to TL specifications. It preserves all semantic details of the prompt to formulate its TL specification. In contrast to previous works that focus solely on formulating relationships between singular objects, isolated events [17, 5, 8], or encoding spatial information [35, 45], our method, PULS, fully integrates both temporal and spatial details which previous methods fall short. By incorporating four evaluation modes, we provide a more comprehensive assessment across spatio-temporal dimensions.
@@
PULS consists of two modules, Text-to-Proposition (T2P) and Text-to-TL (T2TL). Thes use the optimized prompts $\theta^{\star}_{T2P}$ and $\theta^{\star}_{T2TL}$ respectively. Each optimized prompt includes carefully selected few-shot examples from a compiled training dataset $\mathfrak{D}_{\text{T2TL}|\text{train}}$, that maximize the accuracy of having correct $\mathcal{P}$ and $\Phi$ given $m\in M$.
@@ 4.2 Semantic Score from Neural Perception Model
We use the VLM to interpret semantics [2, 43, 34, 54] and extract confidence scores from $\mathcal{F}$ based on the text query $T$ We pass each $p_{i}\in\mathcal{P}$ along with the prompt to the VLM and calculate the token probability for the output response, which is either True or False. To calculate the token probability, we retrieve logits for the response tokens and compute the probability of that token after applying a softmax. Finally, the semantic confidence score is the product of these probabilities as follows:
@@
where $(t_{1},\dots,t_{k})$ is the sequence of tokens in the response. Each term $P(t_{j}|\cdot)=\frac{e^{l_{j,t_{j}}}}{\sum_{z}e^{l_{j,z}}}$ is over the logits $l_{j,{t_{j}}}$ at position $j$, whereas $P(t_{k}|\cdot)=\frac{e^{l_{j,t_{k}}}}{\sum_{z}e^{l_{k,z}}}$ is over those at position $k$. The prompt for the semantic detector VLM can be found in the Appendix.
@@
                        $c_{i}=\mathcal{M}_{\text{VLM}}(p_{i},\mathcal{F})$
@@
                        $C^{\star}[p_{i,n}]\leftarrow f_{\text{VLM}}(c_{i};\gamma_{fp})$
@@
22      $S_{\textit{NeuS-V}}\leftarrow\frac{\sum S}{N}$
23
@@ 4.4 Verifying Synthetic Video Formally
$S_{\textit{NeuS-V}}\leftarrow\frac{\sum S}{N}.$ (10)
@@ 5 Experimental Setup
As we observe that existing benchmarks for T2V evaluation [23, 39, 26] lack explicit temporal instructions, we present the NeuS-V prompt suite with a total of 160 prompts. These are designed to rigorously evaluate models on their ability to maintain temporal coherence and accurately follow event sequences. The suite spans four themes (“Nature”, “Human & Animal Activities”, “Object Interactions”, and “Driving Data”) across three complexity levels (basic, intermediate, and advanced) based on the number of temporal and logical operators.
@@ 6.1 Formal Evaluation of Text-to-Video Models
As outlined in Section 5, we benchmark both closed-source and open-source state-of-the-art models: Gen-3 and Pika (closed-source), alongside T2V-Turbo-v2 [33] and CogVideoX-5B [52] (open-source).
@@
We break down and analyze the performance of each T2V model on our evaluation suite, as shown in Table 1, organized by theme and complexity. Results indicate that most models perform best when generating videos on the themes “Nature” and “Human and Animal Activities”. The Gen-3 model achieves the highest scores across the suite, a result corroborated by human annotators. Furthermore, Figure 5 illustrates how each of the four evaluation modes in PULS are utilized between models. Different prompts rely on distinct modes, indicating that a combination of these modes comprehensively captures the variety of temporal specifications in the prompts.
@@
By Theme Nature 0.630 (0.47) 0.416 (0.70) 0.449 (0.46) 0.486 (0.53)
Human & Animal Activities 0.639 (0.80) 0.418 (0.67) 0.444 (0.66) 0.492 (0.43)
Object Interactions 0.602 (0.16) 0.402 (0.40) 0.398 (0.66) 0.429 (0.65)
Driving Data 0.586 (0.48) 0.399 (0.66) 0.375 (0.30) 0.437 (0.52)
By Complexity Basic (1 TL op.) 0.661 (0.60) 0.475 (0.70) 0.475 (0.58) 0.495 (0.65)
Intermediate (2 TL ops.) 0.555 (0.27) 0.352 (0.44) 0.360 (0.38) 0.415 (0.28)
Advanced (3 TL ops.) 0.605 (-0.01) 0.314 (0.33) 0.349 (0.42) 0.460 (0.78)
Overall Score 0.614 (0.48) 0.409 (0.62) 0.417 (0.55) 0.461 (0.54)
Table 1: Benchmarking SOTA Text-to-Video Models. Performance breakdown with correlations to human evaluation (in brackets).
@@ 7 Discussion
Evaluations in Table 1 reveal a significant limitation in current text-to-video models when handling temporally extended prompts. Although marketed as text-driven video generation, these models often fall short of creating motion and sequences that truly follow the given prompts. Much of the perceived “motion” comes from simplistic effects like panning and zooming with minimal capacity for generating interactions or meaningful scene transitions. In practice, video-generation workflows require numerous rounds of re-prompting to achieve prompt adherence, whereas ideally, we should expect alignment with the prompt most times.
@@ 9 Temporal Logic Operation Example
Given a set of atomic propositions $\mathcal{P}=$Event A, Event B}, the TL specification $\Phi=$Box Event A (read as “Always Event A”) means that ‘Event A’ is True for every step in the sequence. Additionally, $\Phi=$diamondsuit Event B (read as “eventually event b”) indicates that there exists at least one ‘Event B’ in the sequence. Lastly, $\Phi=$ Event A $\mathsf{U}$  Event B (read as “Event A Until Event B”) means that ‘Event A’ exists until ‘Event B’ becomes True, and then ‘Event B’ remains True for all future steps.
@@ 10 Prompt Understanding via Temporal Logic Specification (PULS)
In order to fully parse $\mathfrak{D}_{\text{train}}$, we first begin with the larger dataset $\mathfrak{D}$ with size $\mathcal{B}$ where $\mathfrak{D}=\mathfrak{D}_{\text{T2P}}\cup\mathfrak{D}_{\text{T2TL}}$. $\mathfrak{D}_{\text{T2P}}$ and $\mathfrak{D}_{\text{T2TL}}$ are defined as the following:
@@ 10.1 DSPy & MIPROv2
T2P Prompt for PULS. System prompt to map prompts to the propositions for the “Overall Consistency” mode. Other modes follow a similar structure, albeit with minor changes in instructions and few-shot examples.
@@ 11.1 Inference Via Vision Language Models
We use VLM as a semantic detector as explained in Section 4.2. We pass each atomic proposition $p_{i}\in\mathcal{P}$ such as “person”, “car”, “person in the car”, etc. along with Section 11.1. Once the VLM outputs either ‘Yes’ or ‘No’, we compute the token probability of the response and use it as a confidence score for the detection.
@@
T2TL Prompt for PULS. System prompt to map prompts and propositions for the “Overall Consistency” mode. Other modes follow a similar structure, albeit with minor changes in instructions and few-shot examples.
@@ 13 NeuS-V Prompt Suite
Our dataset is carefully designed to evaluate temporal fidelity and event sequencing in generated videos. It spans four themes – “Nature”, “Human & Animal Activities”, “Object Interactions”, and “Driving Data”, with each theme containing 40 prompts. These prompts vary in complexity, categorized into 20 basic, 15 intermediate, and five advanced prompts based on the number of temporal and logical operators (see Table 5). These prompts were generated using GPT-4o with the system prompt in Section 13. To ensure quality, each prompt was manually verified for clarity, completeness, and relevance.
@@ 38       $\mathcal{A}_{\mathcal{V}}\leftarrow(Q,q_{0},\delta,\lambda)$
Object Existence: (“snow”) $\mathsf{U}$ (“ground”)
Object Action Alignment: (“snow_falls” $\mathsf{U}$ “ground_is_covered”)
Spatial Relationship: F (“snow_covers_ground”)
Overall Consistency: (“snow_falling” $\mathsf{U}$ “it_covers_the_ground”)
@@
Object Existence: G (“waves” & “shore”)
Object Action Alignment: G (“waves_crash_against_rocky_shore”)
Spatial Relationship: G (“waves_on_shore”)
Overall Consistency: G (“waves_crashing_against_the_rocky_shore”)
@@
Object Existence: (“sun_shining” $\mathsf{U}$ “clouds_gather”) & F (“rain_begins_to_fall”)
Object Action Alignment: (“sun_shines” $\mathsf{U}$ “clouds_gather”) & F (“rain_falls”)
Spatial Relationship: G (“sun_over_horizon” & “dew_on_grass”)
Overall Consistency: (“sun_shining” $\mathsf{U}$ “clouds_gather”) & F (“rain_begins_to_fall”)
@@
Object Existence: (“butterfly” & “flower”) $\mathsf{U}$ (“wind”)
Object Action Alignment: (“butterfly_rests_on_flower” $\mathsf{U}$ “gust_of_wind_comes”) & F (“butterfly_flies_away”)
Spatial Relationship: (“butterfly_on_flower”) $\mathsf{U}$ (“butterfly_flies_away”)
Overall Consistency: (“butterfly_resting_on_a_flower” $\mathsf{U}$ “gust_of_wind_comes”) & F (“it_flies_away”)
@@
Object Existence: G (“river” & “storm_clouds”)
Object Action Alignment: ((“river_flows_quietly” $\mathsf{U}$ “sky_darkens”)) & F(“rain_pours”)
Spatial Relationship: G(“river_flowing_through_valley”) $\mathsf{U}$ (“sky_darkens_with_storm_clouds”) & F “rain_begins_to_pour_heavily”
Overall Consistency: G ((“river_flowing_quietly_through_the_valley”) $\mathsf{U}$ “sky_darkens_with_storm_clouds”) & F “rain_begins_to_pour_heavily”
@@
Object Existence: (“dog”) $\mathsf{U}$ (“ball”)
Object Action Alignment: (“dog_barks” $\mathsf{U}$ “someone_throws_ball”)
Spatial Relationship: G (“clock_on_mantle” & “fireplace_beneath_mantle”)
Overall Consistency: (“dog_barking” $\mathsf{U}$ “someone_throws_a_ball”)
@@
Object Existence: (“bird”) $\mathsf{U}$ (“branch”)
Object Action Alignment: (“bird_sings” $\mathsf{U}$ “bird_flies_away_to_another_branch”)
Spatial Relationship: (“bird_on_branch”) $\mathsf{U}$ (“bird_on_different_branch”)
Overall Consistency: (“bird_singing” $\mathsf{U}$ “it_flies_away_to_another_branch”)
@@
Object Existence: (“child” & “sandcastle”) $\mathsf{U}$ “tide”
Object Action Alignment: ((“child_builds_sandcastle” $\mathsf{U}$ “tide_rises”)) & F(“child_watches_it_wash_away”)
Spatial Relationship: (“child_building_sandcastle” $\mathsf{U}$ “tide_rises”)
Overall Consistency: (“child_building_a_sandcastle” $\mathsf{U}$ “tide_rises”) & F “child_watches_sandcastle_wash_away”
@@
Object Existence: G (“cat” & “butterflies”)
Object Action Alignment: G((“cat_lounges” & “butterflies_flutter”))
Spatial Relationship: G(“cat_lounging_on_porch” & “butterflies_fluttering_around”)
Overall Consistency: G (“cat_lounging_on_the_porch” & “butterflies_fluttering_around”)
@@
Object Existence: (“dog” & “backyard”) $\mathsf{U}$ (“owner” & “ball”)
Object Action Alignment: (“dog_digs_in_backyard” $\mathsf{U}$ “owner_arrives”) & F (“dog_plays_fetch_with_owner”)
Spatial Relationship: (“dog_in_backyard”) $\mathsf{U}$ (“owner_arrives”)
Overall Consistency: (“dog_digging_in_the_backyard” $\mathsf{U}$ “its_owner_arrives”) & F (“they_play_fetch_together”)
Table 6: NeuS-V Prompt Suite: Illustrative prompts and their detailed specifications (across all four modes) for varying complexities within the “Nature” and “Human and Animal Activities” themes.
@@
Object Existence: (“lamp”) $\mathsf{U}$ (¡‘lamp”)
Object Action Alignment: (“lamp_glows” $\mathsf{U}$ “lamp_is_turned_off”)
Spatial Relationship: (“cars_passing_by_person”)
Overall Consistency: (“lamp_glowing” $\mathsf{U}$ “it_is_turned_off”)
@@
Object Existence: (“car_engine” & “dashboard_lights”)
Object Action Alignment: (“car_engine_runs”&“dashboard_flashes”)
Spatial Relationship: (“car_engine_running” & “dashboard_lights_flashing”)
Overall Consistency: (“car_engine_running” & “dashboard_lights_flashing”)

### v5

@@ Abstract
Recent advancements in text-to-video models such as Sora, Gen-3, MovieGen, and CogVideoX are pushing the boundaries of synthetic video generation, with adoption seen in fields like robotics, autonomous driving, and entertainment. As these models become prevalent, various metrics and benchmarks have emerged to evaluate the quality of the generated videos. However, these metrics emphasize visual quality and smoothness, neglecting temporal fidelity and text-to-video alignment, which are crucial for safety-critical applications. To address this gap, we introduce NeuS-V, a novel synthetic video evaluation metric that rigorously assesses text-to-video alignment using neuro-symbolic formal verification techniques. Our approach first converts the prompt into a formally defined Temporal Logic (TL) specification and translates the generated video into an automaton representation. Then, it evaluates the text-to-video alignment by formally checking the video automaton against the TL specification. Furthermore, we present a dataset of temporally extended prompts to evaluate state-of-the-art video generation models against our benchmark. We find that NeuS-V demonstrates a higher correlation by over 5×5$\times$5 × with human evaluations when compared to existing metrics. Our evaluation further reveals that current video generation models perform poorly on these temporally complex prompts, highlighting the need for future work in improving text-to-video generation capabilities. We open-source our benchmark, code, and dataset at utaustin-swarmlab.github.io/NeuS-V.
@@ 1 Introduction
Various evaluation metrics have emerged to identify well-aligned synthetic videos to the prompt from recently developed T2V models [20, 14, 4, 57, 33]. While many video evaluation tools [41, 39, 23] focus on visual quality, some approaches use vision-language models [50, 51] or video captioning models [9] to evaluate the semantic alignment between the generated video and the original prompt. More notably, VBench [23] evaluates videos across multiple dimensions and categories and is recently being used in leaderboards***https://huggingface.co/spaces/Vchitect/VBench_Leaderboard to evaluate T2V models. However, naive evaluation of a video generated by one neural network using another is neither rigorous nor interpretable. Moreover, current evaluation metrics still miss critical spatio-temporal aspects that are intrinsic to video data (See Fig. 1).
@@ 3 Preliminaries
In the following sections, we present a running example to help illustrate our approach. Let’s say we generate an autonomous driving video using T2V models with the prompt – “A car drives down a road on a clear sunny day, interacting with cyclists who signal turns to avoid obstacles".
@@
$\begin{array}[]{rl}\mathcal{P}&=\{\text{car driving},\ \text{clear day},\ \text{cyclist signals turn},\\ &\quad\ \ \text{cyclist turns},\ \text{cyclist avoids obstacle}\},\\ \Phi&=\Box\Big{(}(\text{car driving}\wedge\text{clear day})\wedge(\text{cyclist signals turn})\end{array}$ (1)
@@ 4.1 Prompt Understanding via Temporal Logic Specification
We propose Prompt Understanding via temporal Logic (PULS), a novel method that facilitates an efficient and accurate conversion of text prompts to TL specifications. It preserves all semantic details of the prompt to formulate its TL specification. In contrast to previous works that focus solely on formulating relationships between singular objects, isolated events [17, 5, 8], or encoding spatial information [37, 47], our method, PULS, fully captures both temporal and spatial details. By incorporating four evaluation modes, we provide a more comprehensive assessment across spatio-temporal dimensions.
@@
PULS consists of two modules, Text-to-Proposition (T2P) and Text-to-TL (T2TL). These use the optimized prompts $\theta^{\star}_{T2P}$ and $\theta^{\star}_{T2TL}$ respectively. Each optimized prompt includes carefully selected few-shot examples from a compiled training dataset $\mathfrak{D}_{\text{T2TL}|\text{train}}$, that maximize the accuracy of having correct $\mathcal{P}$ and $\Phi$ given $m\in M$.
@@ 4.2 Semantic Score from Neural Perception Model
We use the VLM to interpret semantics [2, 45, 36, 56] and extract confidence scores from $\mathcal{F}$ based on the text query $T$ We pass each $p_{i}\in\mathcal{P}$ along with the prompt to the VLM and calculate the token probability for the output response, which is either True or False. To calculate the token probability, we retrieve logits for the response tokens and compute the probability of that token after applying a softmax. Next, the semantic confidence score is the product of these probabilities as follows:
@@
where $(t_{1},\dots,t_{k})$ is the sequence of tokens in the response. The probability of the token at $j$’th response token is $P(t_{j}|\cdot)=\frac{e^{l_{j,t_{j}}}}{\sum_{z}e^{l_{j,z}}}$, where $l_{j,*}$ is the logit vector, and $l_{j,{t_{j}}}$ is the logit corresponding to token $t_{j}$. Finally, we calibrate the VLM to map the confidence to the probability using the semantic score mapping function $c^{\star}_{i}\leftarrow f_{\text{VLM}}(c_{i};\gamma_{fp})$, where $c^{\star}_{i}$ is the calibrated confidence and $\gamma_{fp}$ is the false positive threshold. We provide additional details about the calibration process and the prompt design for the semantic detector VLM in the Appendix.
@@
                        $c_{i}\leftarrow\mathcal{M}_{\text{VLM}}(p_{i},\mathcal{F})$
@@
                        $c^{\star}_{i}\leftarrow f_{\text{VLM}}(c_{i};\gamma_{fp})$
@@
                        $C^{\star}[p_{i,n}]\leftarrow c^{\star}_{i}$
                          // Append $c^{\star}_{i}$ to the semantic confidence set
16
@@
23      $S_{\textit{NeuS-V}}\leftarrow\frac{\sum S}{\mid S\mid}$
24
@@ 4.4 Verifying Synthetic Video Formally
$S_{\textit{NeuS-V}}\leftarrow\frac{\sum S}{\mid S\mid}.$ (10)
@@ 5 Experimental Setup
As we observe that existing benchmarks for T2V evaluation [23, 41, 26] lack explicit temporal instructions, we present the NeuS-V prompt suite with a total of 360 prompts. These are designed to rigorously evaluate models on their ability to maintain temporal coherence and accurately follow event sequences. The suite spans four themes (“Nature", “Human & Animal Activities", “Object Interactions", and “Driving Data") across three complexity levels (basic, intermediate, and advanced) based on the number of temporal and logical operators.
@@ 6.1 Formal Evaluation of Text-to-Video Models
As outlined in Section 5, we benchmark both closed-source and open-source state-of-the-art models: Gen-3 and Pika (closed-source), alongside T2V-Turbo-v2 [34, 35] and CogVideoX-5B [54] (open-source).
@@
We break down and analyze the performance of each T2V model on our evaluation suite, as shown in Table 1, organized by theme and complexity for a subset of 160 prompts. Results indicate that most models perform best when generating videos on the themes “Nature" and “Human and Animal Activities". The Gen-3 model achieves the highest scores across the suite, a result corroborated by human annotators. Furthermore, Figure 5 illustrates how each of the four evaluation modes in PULS are utilized between models. Different prompts rely on distinct modes, indicating that a combination of these modes comprehensively captures the variety of temporal specifications in the prompts.
@@
By Theme Nature 0.716 (0.47) 0.479 (0.70) 0.564 (0.46) 0.580 (0.53)
Human & Animal Activities 0.752 (0.80) 0.531 (0.67) 0.564 (0.66) 0.623 (0.43)
Object Interactions 0.710 (0.16) 0.500 (0.40) 0.553 (0.66) 0.573 (0.65)
Driving Data 0.716 (0.48) 0.525 (0.66) 0.525 (0.30) 0.580 (0.52)
By Complexity Basic (1 TL op.) 0.774 (0.60) 0.589 (0.70) 0.610 (0.58) 0.641 (0.65)
Intermediate (2 TL ops.) 0.680 (0.27) 0.464 (0.44) 0.508 (0.38) 0.549 (0.28)
Advanced (3 TL ops.) 0.692 (-0.01) 0.400 (0.33) 0.494 (0.42) 0.550 (0.78)
Overall Score 0.723 (0.48) 0.508 (0.62) 0.552 (0.55) 0.589 (0.54)
Table 1: Benchmarking SOTA Text-to-Video Models. Performance metrics reflect the full 360-prompt set, while correlations to human evaluations (in parentheses) are computed on a 160-prompt subset. NeuS-V enjoys high correlation across all themes and complexities.
@@ 7 Discussion
Evaluations in Table 1 reveal a significant limitation in current text-to-video models when handling temporally extended prompts. Although marketed as text-driven video generation, these models often fall short of creating motion and sequences that truly follow the given prompts. Much of the perceived “motion" comes from simplistic effects like panning and zooming with minimal capacity for generating interactions or meaningful scene transitions. In practice, video-generation workflows require numerous rounds of re-prompting to achieve prompt adherence, whereas ideally, we should expect alignment with the prompt most times.
@@ 8 Conclusion
Acknowledgments
This material is based upon work supported in part by the Office of Naval Research (ONR) under Grant No. N00014-22-1-2254. Additionally, this work was supported by the Defense Advanced Research Projects Agency (DARPA) contract DARPA ANSR: RTX CW2231110. Approved for Public Release, Distribution Unlimited.
Appendix
@@ 9 Temporal Logic Operation Example
Given a set of atomic propositions $\mathcal{P}=$Event A, Event B}, the TL specification $\Phi=$Box Event A (read as “Always Event A") means that ‘Event A’ is True for every step in the sequence. Additionally, $\Phi=$diamondsuit Event B (read as “eventually event b") indicates that there exists at least one ‘Event B’ in the sequence. Lastly, $\Phi=$ Event A $\mathsf{U}$  Event B (read as “Event A Until Event B") means that ‘Event A’ exists until ‘Event B’ becomes True, and then ‘Event B’ remains True for all future steps.
@@ 10 Prompt Understanding via Temporal Logic Specification (PULS)
In order to obtain $\mathfrak{D}_{\text{train}}$, we first begin with the larger dataset $\mathfrak{D}$ with size $\mathcal{B}$ where $\mathfrak{D}=\mathfrak{D}_{\text{T2P}}\cup\mathfrak{D}_{\text{T2TL}}$. $\mathfrak{D}_{\text{T2P}}$ and $\mathfrak{D}_{\text{T2TL}}$ are defined as the following:
@@ 10.1 DSPy & MIPROv2
T2P Prompt for PULS. System prompt to map prompts to the propositions for the “Overall Consistency" mode. Other modes follow a similar structure, albeit with minor changes in instructions and few-shot examples.
@@ 11.1 Inference Via Vision Language Models
We use VLM as a semantic detector as explained in Section 4.2. We pass each atomic proposition $p_{i}\in\mathcal{P}$ such as “person", “car", “person in the car", etc. along with Section 11.1. Once the VLM outputs either ‘Yes’ or ‘No’, we compute the token probability of the response and use it as a confidence score for the detection.
@@
T2TL Prompt for PULS. System prompt to map prompts and propositions for the “Overall Consistency" mode. Other modes follow a similar structure, albeit with minor changes in instructions and few-shot examples.
@@ 13 NeuS-V Prompt Suite
Our dataset is carefully designed to evaluate temporal fidelity and event sequencing in generated videos. It spans four themes – “Nature", “Human & Animal Activities", “Object Interactions", and “Driving Data", with each theme containing 40 prompts. These prompts vary in complexity, categorized into 20 basic, 15 intermediate, and five advanced prompts based on the number of temporal and logical operators (see Table 5). These prompts were generated using GPT-4o with the system prompt in Section 13. To ensure quality, each prompt was manually verified for clarity, completeness, and relevance.
@@ 38       $\mathcal{A}_{\mathcal{V}}\leftarrow(Q,q_{0},\delta,\lambda)$
Object Existence: (“snow") $\mathsf{U}$ (“ground")
Object Action Alignment: (“snow_falls" $\mathsf{U}$ “ground_is_covered")
Spatial Relationship: F (“snow_covers_ground")
Overall Consistency: (“snow_falling" $\mathsf{U}$ “it_covers_the_ground")
@@
Object Existence: G (“waves" & “shore")
Object Action Alignment: G (“waves_crash_against_rocky_shore")
Spatial Relationship: G (“waves_on_shore")
Overall Consistency: G (“waves_crashing_against_the_rocky_shore")
@@
Object Existence: (“sun_shining" $\mathsf{U}$ “clouds_gather") & F (“rain_begins_to_fall")
Object Action Alignment: (“sun_shines" $\mathsf{U}$ “clouds_gather") & F (“rain_falls")
Spatial Relationship: G (“sun_over_horizon" & “dew_on_grass")
Overall Consistency: (“sun_shining" $\mathsf{U}$ “clouds_gather") & F (“rain_begins_to_fall")
@@
Object Existence: (“butterfly" & “flower") $\mathsf{U}$ (“wind")
Object Action Alignment: (“butterfly_rests_on_flower" $\mathsf{U}$ “gust_of_wind_comes") & F (“butterfly_flies_away")
Spatial Relationship: (“butterfly_on_flower") $\mathsf{U}$ (“butterfly_flies_away")
Overall Consistency: (“butterfly_resting_on_a_flower" $\mathsf{U}$ “gust_of_wind_comes") & F (“it_flies_away")
@@
Object Existence: G (“river" & “storm_clouds")
Object Action Alignment: ((“river_flows_quietly" $\mathsf{U}$ “sky_darkens")) & F(“rain_pours")
Spatial Relationship: G(“river_flowing_through_valley") $\mathsf{U}$ (“sky_darkens_with_storm_clouds") & F “rain_begins_to_pour_heavily"
Overall Consistency: G ((“river_flowing_quietly_through_the_valley") $\mathsf{U}$ “sky_darkens_with_storm_clouds") & F “rain_begins_to_pour_heavily"
@@
Object Existence: (“dog") $\mathsf{U}$ (“ball")
Object Action Alignment: (“dog_barks" $\mathsf{U}$ “someone_throws_ball")
Spatial Relationship: G (“clock_on_mantle" & “fireplace_beneath_mantle")
Overall Consistency: (“dog_barking" $\mathsf{U}$ “someone_throws_a_ball")
@@
Object Existence: (“bird") $\mathsf{U}$ (“branch")
Object Action Alignment: (“bird_sings" $\mathsf{U}$ “bird_flies_away_to_another_branch")
Spatial Relationship: (“bird_on_branch") $\mathsf{U}$ (“bird_on_different_branch")
Overall Consistency: (“bird_singing" $\mathsf{U}$ “it_flies_away_to_another_branch")
@@
Object Existence: (“child" & “sandcastle") $\mathsf{U}$ “tide"
Object Action Alignment: ((“child_builds_sandcastle" $\mathsf{U}$ “tide_rises")) & F(“child_watches_it_wash_away")
Spatial Relationship: (“child_building_sandcastle" $\mathsf{U}$ “tide_rises")
Overall Consistency: (“child_building_a_sandcastle" $\mathsf{U}$ “tide_rises") & F “child_watches_sandcastle_wash_away"
@@
Object Existence: G (“cat" & “butterflies")
Object Action Alignment: G((“cat_lounges" & “butterflies_flutter"))
Spatial Relationship: G(“cat_lounging_on_porch" & “butterflies_fluttering_around")
Overall Consistency: G (“cat_lounging_on_the_porch" & “butterflies_fluttering_around")
@@
Object Existence: (“dog" & “backyard") $\mathsf{U}$ (“owner" & “ball")
Object Action Alignment: (“dog_digs_in_backyard" $\mathsf{U}$ “owner_arrives") & F (“dog_plays_fetch_with_owner")
Spatial Relationship: (“dog_in_backyard") $\mathsf{U}$ (“owner_arrives")
Overall Consistency: (“dog_digging_in_the_backyard" $\mathsf{U}$ “its_owner_arrives") & F (“they_play_fetch_together")
Table 6: NeuS-V Prompt Suite: Illustrative prompts and their detailed specifications (across all four modes) for varying complexities within the “Nature" and “Human and Animal Activities" themes.
@@
Object Existence: (“lamp") $\mathsf{U}$ (¡‘lamp")
Object Action Alignment: (“lamp_glows" $\mathsf{U}$ “lamp_is_turned_off")
Spatial Relationship: (“cars_passing_by_person")
Overall Consistency: (“lamp_glowing" $\mathsf{U}$ “it_is_turned_off")
@@
Object Existence: (“car_engine" & “dashboard_lights")
Object Action Alignment: (“car_engine_runs"&“dashboard_flashes")
Spatial Relationship: (“car_engine_running" & “dashboard_lights_flashing")
Overall Consistency: (“car_engine_running" & “dashboard_lights_flashing")